Suppose the demand curve is described by the equation Q = 400 - 10P, where Q is the quantity
demanded and P is the price of the good. At price P = 30, the quantity demanded is
a. 400
b. 300
c. 200
d. 100
1
Expert's answer
2020-03-03T09:45:48-0500

Answer

D. 100

Explanation

Q= 400- 10P

To determine the quantity demanded given, price P, we need to replace P in the question.

So in the equation,

Q=400- 10P, replace P which is 30

Q= 400- 10(30)

 =400- 300

 =100

Thus, the quantity demanded is 100.